Shift summary for owl shift 8/17/99

S. Elizabeth Turner (eturner@kudzu)
Tue, 17 Aug 1999 06:46:58 -0500

Shift Summary for owl shift, 1999, August 17.

On Shift: YBH, SET

Hours of beam available at all intensities: 6h 50m
Hours of beam available at nominal intensity: 6h 50m
Hours of E832 data (special runs not included): 6h 35m
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: tape change, a few kfrends
========================================================================
Runs ended this shift:
Run 14241 23.9M events - BEAM_832

Tapes brought to FCC this shift:
(including tapes for online split)
QK*276-278, QKB097
========================================================================
Accesses during shift:
none
========================================================================
Overview of shift:
Smooth E832 datataking
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14241 249 1:26 ktevana errors -- kfrend
321 2:50 ktevana errors -- kfrend
339 3:20 no beam
367 3:55 run ends, tapes full

4:00 mount 3 onsplit tapes

14242 4:00 change tapes, init run
38 5:00 DAQDEAD_SEC = 1.47
39 lose beam
43 beam returns
91 no beam
========================================================================
Follow up Notes on Previous Shift

Tape QKK276 is missing its box -- the box for QKK275 (which was
vaulted last shift) is here -- so we suspect a swap. We will ask
FCC to unvault QKK275, then swap boxes, then vault everything.
========================================================================
Items to Follow up on Next Shift

Monitor CSI comparator problem for channel 961 -- right now the
problem occurs at low frequency.